# Exploring Multilingual Human Value Concepts in Large Language Models: Is Value Alignment Consistent, Transferable and Controllable across Languages?

### v1

@@ 3.4 Recognizing Cross-Lingual Concepts
To investigate the cross-lingual transferability of a specific concept of human value across different languages, we propose a method for cross-lingual concept recognition. Given two languages, $l_{1}$ and $l_{2}$, we calculate how accurately $\bm{v}_{c}^{l_{1}}$ and $\bm{v}_{c}^{l_{2}}$ can be used to recognize the concept $c$ in language $l_{2}$, resulting in $\textnormal{Acc}_{c}^{l_{1}\rightarrow l_{2}}$ and $\textnormal{Acc}_{c}^{l_{2}}$. The inequality $\textnormal{Acc}_{c}^{l_{1}\rightarrow l_{2}}\geq\textnormal{Acc}_{c}^{l_{2}}$ being true signifies the successful transfer of concept $c$ from $l_{1}$ to $l_{2}$. Conversely, we calculate $\textnormal{Acc}_{c}^{l_{2}\rightarrow l_{1}}$ and $\textnormal{Acc}_{c}^{l_{1}}$ to explore the transferability of concept $c$ from $l_{2}$ to $l_{1}$.
Figure 2: Multilingual concept recognition accuracy of LLaMA2-chat, Qwen-chat and BLOOMZ series, averaged across all human values. The performance of the three 7B-sized models are connected with dashed lines for performance comparison.
Figure 3: English concept recognition accuary with varying numbers of training samples for collecting concept vectors. The result are based on LLaMA2-chat-13B. We calculate the average accuracy across all layers to ensure the results of different settings are comparable.
@@ 4 Experiments
We conducted extensive experiments with the proposed framework on 7 human values, 16 languages and 3 LLM families to answer questions Q1, Q2 and Q3. We leave the question Q4 to Section 5.
@@ 4.1 Experimental Setup
We conducted experiments on the following human values: morality, deontology, utilitarianism, fairness, truthfulness, toxicity and harmfulness. We utilized three subsets of the ETHICS dataset (Hendrycks et al., 2021) for morality, deontology, and utilitarianism. Regarding fairness, truthfulness, toxicity, and harmfulness, we chose the StereoSet (Nadeem et al., 2021), TruthfulQA (Lin et al., 2022), REALTOXICITYPROMPTS (Gehman et al., 2020), AdvBench (Zou et al., 2023b) dataset, respectively. Please refer to Appendix A for detailed definitions, data splits, and examples of each human value.
@@
We translated the aforementioned human value datasets from English into 15 non-English languages using Google Translate. These languages belong to various language families, including Indo-European (Catalan, French, Indonesian, Portuguese, Spanish), Niger-Congo (Chichewa, Swahili), Dravidian (Tamil, Telugu), Uralic (Finnish, Hungarian), Sino-Tibetan (Chinese), Japonic (Japanese), Koreanic (Korean) and Austro-Asiatic (Vietnamese).
@@
Our experiments involved three multilingual LLM families, including the LLaMA2-chat series (7B, 13B, 70B) (Touvron et al., 2023), Qwen-chat series (1B8, 7B, 14B) (Bai et al., 2023) and BLOOMZ series (560M, 1B7, 7B1) (Scao et al., 2022). Appendix B provides detailed language distributions of their pre-training data.
@@
4.2 Do LLMs Encode Concepts Representing Human Values in Multiple Languages?
@@
Figure 2 illustrates the multilingual concept recognition accuracy of the three LLM families, averaged across all human values. [^4] See Appendix C for complete results of each human value and extra discussions. We observe that all three models achieve notable accuracy across all represented languages [^5] We also observe that the performance in unrepresented languages consistently surpasses the random baseline. The model’s understanding in these languages may stem from cross-lingual transfer from other languages. Qwen’s technical report only mentions the inclusion of English and Chinese in its pre-training data. We conjecture the inclusion of 10 other languages (fr,es,pt,vi,ca,id,ja,ko,fi,hu) based on its significant recognition performance in these languages. and even the smallest models surpass $\tau=65\%$ accuracy in them. These results demonstrate that LLMs effectively encode human values in a multilingual context.
Figure 2 also shows a clear pattern that increasing model size substantially improves concept distinguishing accuracy, indicating that large models more explicitly encode multilingual human values than small models.
@@
Figure 4: Cross-lingual similarity of concept vectors across all language pairs, averaged over all human values. The languages included in each model’s pre-training data are presented and sorted based on their proportions in the corresponding model’s pre-training data. For Qwen-chat series, we conjecture its language inclusion based on multilingual concept recognition accuracy (Section 4.2) and display its primary languages, zh and en, at the forefront.
4.2.1 Varying the Size of $\mathcal{T}_{c}^{\textnormal{train}}$
We employed varying amounts of training samples to extract concept vectors, and the recognition performance for each human value is illustrated in Figure 3. Surprisingly, optimal accuracy can be achieved for all human values even with few training samples, consistent with the findings by Li et al. (2023), suggesting that the concept vectors for human values are readily extractable in LLMs.
Furthermore, we observe notable differences in the recognition accuracy of different human values, indicating different degrees of difficulty in capturing them. Specifically, harmfulness, toxicity, morality, and deontology are relatively explicitly encoded human values. In contrast, LLMs encounter a greater challenge in recognizing concepts like truthfulness, fairness and utilitarianism.
@@ 4.2.1 Varying the Size of $\mathcal{T}_{c}^{\textnormal{train}}$
4.3 To What Extent are Human Value Concepts Consistent and Transferable across Different Languages?
@@
Through computing cross-lingual similarity of concept vectors (§3.3) and recognizing cross-lingual concepts [^6] If not otherwise specified, concepts in the experiments refer to the 7 types of human value concepts.3.4), we investigated the cross-lingual consistency and transferability of these human value concepts (Q2). Moreover, analyzing these concepts on LLMs trained with different multilingual data distributions provides insights into the multilinguality of LLMs (Q3).
@@ 4.3.1 Trait 1: Inconsistency of Concept Representations between High- and Low-Resource Languages
Figure 4 illustrates the cross-lingual similarity of concept vectors captured by the three 7B-sized models. [^7] Similarity results across all model sizes and extra discussions are detailed in Appendix E. We find that different multilinguality leads to different patterns of cross-lingual concept consistency. In the case of LLaMA2-chat-7B, the absolute dominance of English results in the model learning relatively independent concept representations for English, showing concept representation inconsistency between English and other languages, while higher cross-lingual concept consistency is observed among other languages. BLOOMZ-7B1’s cross-lingual concept consistency exhibits a very different pattern: the four languages with the lowest proportions (ta, te, sw, ny, accounting for 0.50%, 0.19%, 0.015%, and 0.00007% of pre-training data, respectively) show the lowest concept consistency (similarity) with other languages, while languages with relatively higher proportions (en with the highest percentage of 30.04%, and ca with the lowest percentage of 1.10%) demonstrate higher concept consistency with each other. [^8] We observe inconsistency between Spanish and other languages in BLOOMZ-7B1. We would like to explore this in our future work. For Qwen-chat-7B, we do not observe significant cross-lingual consistency between the main languages (zh, en) and other languages. In summary, cross-lingual concept inconsistency is more likely to occur between high- and low-resource languages.
@@
Direct Category Direct Category Direct Category Direct Category
@@
Table 1: Pearson correlation between cross-lingual concept consistency and linguistic similarity for all language pairs. “Direct” refers to results obtained through direct computation; “Category” pertains to the average results derived by first categorizing languages based on language resources and then computing correlations within different language categories.
Figure 5: Cross-lingual concept transferrability across all language pairs, averaged over all human values. Languages are sorted based on their percentages in the pre-training data.
@@ 4.3.2 Trait 2: Linguistic Relationships Distortion due to the Imbalance of Language Data
We employed two calculation methods to estimate the correlation. The first method directly computes the Pearson correlation on all language pairs (Direct), while the second starts by categorizing language pairs based on language resources. Subsequently, correlations are computed within different categories and averaged (Category). Please refer to Appendix D for details of the latter method.
@@
In terms of linguistic variations, cross-lingual concept consistency exhibits the strongest correlation with genetic and syntactic similarity. In contrast, there is a weak positive correlation between cross-lingual concept consistency with geographic similarity, while no correlation is observed with phonological similarity. The results suggest that LLMs embed more consistent concepts of human values for language pairs with similar syntactic structures, genetic relations, and geographic proximity, aligning with previous findings on multilingual factual knowledge (Qi et al., 2023).
@@ 4.3.3 Trait 3: Unidirectional Concept Transfer from High- to Low-Resource Languages
For a given source language $l_{1}$ and target language $l_{2}$, we compute $\textnormal{Acc}_{c}^{l_{1}\rightarrow l_{2}}-\textnormal{Acc}_{c}^{l_{2}}$ (the difference in accuracy scores) to measure the transferability of concept $c$ from $l_{1}$ to $l_{2}$ (Section 3.4). We average differences in accuracy scores over all human values to measure the overall transferability. If the average difference is greater than 0, it indicates positive transferability from $l_{1}$ to $l_{2}$.
We present the cross-lingual concept transferability of the three 7B-sized models in Figure 5.101010Cross-lingual concept transferability across all model sizes and additional discussions are detailed in Appendix F. It provides insights into the influence of LLMs’ multilinguality. Firstly, based on the results of LLaMA- and Qwen-chat-7B, we observe a monotonic concept transfer pattern introduced by the presence of dominant languages. This pattern is characterized by a unidirectional transfer from the dominant language to other languages. This pattern also exhibits an upper triangular cross-lingual transferability (the dashed triangular in Figure 5), indicating that cross-lingual concept transfer from high- to low-resource languages is more prevalent. In contrast, BLOOMZ-7B1 exhibits a relatively balanced bidirectional cross-lingual concept transferability, while for languages with extremely low resources, the tendency of unidirectional transfer persists.
@@
LLaMA2 -chat-7B No-Control 0.97 1.94 6.8 1.94 6.8 4.85 8.74 5.83 3.88 10.68 14.56 4.85 6.44
@@
LLaMA2 -chat-13B No-Control 0.97 0.97 5.83 1.94 5.83 5.83 27.18 8.74 2.91 10.68 15.53 6.8 8.38
@@
En-Control 88.35 99.03 95.15 98.06 97.09 98.06 93.2 94.17 99.03 97.09 90.29 87.38 95.32
LLaMA2 -chat-70B No-Control 0.0 1.94 4.85 0.97 6.8 2.91 27.18 11.65 2.91 20.39 18.45 10.68 9.89
@@
5 Is Value Alignment of LLMs Controllable across Languages?
LLaMA2-chat models, trained with alignment techniques such as RLHF, exhibit value alignment capabilities like rejecting harmful instructions. In this section, we employed the representation engineering (RE) methodology (Zou et al., 2023a) to bypass such defense and further explored the potential for cross-lingual control of value alignment.
@@ 5.1 Cross-Lingual Value Alignment Control
To control a LLM to exhibit behavior aligned with the concept of a human value $c$, a straightforward RE-style method is multiplying the previously extracted concept vector $\bm{v}_{c}$ by a control strength $s$ and adding it to the hidden states of multiple layers $L$ within the target model. This procedure is iteratively applied to each token, formulated as $\bm{h_{i}}^{{}^{\prime}}=\bm{h_{i}}+s\cdot\bm{v}_{c}$, where $\bm{h_{i}}$ and $\bm{h_{i}}^{{}^{\prime}}$ denote the original and perturbed hidden state of $i$-th token, respectively. [^11] Reflecting on Section 3.1, each layer has its specific concept vector, and the perturbation is executed across multiple layers $L$. We omit the detail here for simplicity. In a cross-lingual scenario, we leverage the concept vector $\bm{v}_{c}^{l}$ of the source language ${l}$ to control the model’s behavior across various target languages. To determine appropriate control strength $s$ and control layers $L$ for cross-lingual control, we first conduct hyperparameter search to choose the combination that demonstrates the most effective control on language ${l}$. Subsequently, we employ this combination for cross-lingual control across all target languages and evaluate the control effect on each of them.
In our experiments, a successful control is steering the LLM to follow a harmful instruction rather than rejecting it. We compute the Following rate, representing the proportion of harmful instructions the model follows, to assess the effectiveness of model control. Specifically, we utilize the multilingual negative testing data (harmful instructions) for harmfulness concept (Section 4.1), calculating the Following rate in each language.
Please refer to Appendix G for details of hyperparameter search and model control evaluation.

### v2

@@ 3.4 Recognizing Cross-Lingual Concepts
To investigate the cross-lingual transferability of a specific value concept across languages, we propose a method for cross-lingual concept recognition. Given two languages, $l_{1}$ and $l_{2}$, we calculate how accurately $\bm{v}_{c}^{l_{1}}$ and $\bm{v}_{c}^{l_{2}}$ can be used to recognize the concept $c$ in language $l_{2}$, resulting in $\textnormal{Acc}_{c}^{l_{1}\rightarrow l_{2}}$ and $\textnormal{Acc}_{c}^{l_{2}}$. The inequality $\textnormal{Acc}_{c}^{l_{1}\rightarrow l_{2}}\geq\textnormal{Acc}_{c}^{l_{2}}$ being true signifies the successful transfer of concept $c$ from $l_{1}$ to $l_{2}$. Conversely, we calculate $\textnormal{Acc}_{c}^{l_{2}\rightarrow l_{1}}$ and $\textnormal{Acc}_{c}^{l_{1}}$ to explore the transferability of concept $c$ from $l_{2}$ to $l_{1}$. While evaluating transferability based solely on accuracy changes might imply a unidirectional transfer from high- to low-performing languages, Appendix H.1 indicates that transferability is not solely determined by language performance.
@@ 4 Experiments
We conducted extensive experiments with the proposed framework on 7 human values, 16 languages and 3 LLM families to answer questions Q1, Q2 and Q3. We leave the question Q4 to §5.
@@ 4.1 Experimental Setup
We explored the following values: commonsense morality, deontology, utilitarianism, fairness, truthfulness, toxicity and harmfulness. We utilized 3 subsets of ETHICS dataset (Hendrycks et al., 2021) for commonsense morality, deontology, and utilitarianism. Regarding fairness, truthfulness, toxicity, and harmfulness, we chose the StereoSet (Nadeem et al., 2021), TruthfulQA (Lin et al., 2022), REALTOXICITYPROMPTS (Gehman et al., 2020), AdvBench (Zou et al., 2023b) dataset, respectively.
Appendix B details the sources, data splits, and positive and negative examples for each value.
@@
We translated the aforementioned human value datasets from English into 15 non-English languages using Google Translate. These languages belong to various language families, including Indo-European (Catalan, French, Indonesian, Portuguese, Spanish), Niger-Congo (Chichewa, Swahili), Dravidian (Tamil, Telugu), Uralic (Finnish, Hungarian), Sino-Tibetan (Chinese), Japonic (Japanese), Koreanic (Korean) and Austro-Asiatic (Vietnamese). The impact of translation quality on our results is discussed in Appendix C.
@@
Our experiments involved three multilingual LLM families, including the LLaMA2-chat series (7B, 13B, 70B) (Touvron et al., 2023), Qwen-chat series (1B8, 7B, 14B) (Bai et al., 2023) and BLOOMZ series (560M, 1B7, 7B1) (Scao et al., 2022). Appendix D provides detailed language distributions of their pre-training data.
@@
4.2 Q1: Do LLMs Encode Concepts Representing Human Values in Multiple Languages?
@@
Figure 1 illustrates the multilingual concept recognition accuracy of the three LLM families, averaged across all value concepts. We observe that all three models achieve notable accuracy across all represented languages [^2] The model’s understanding in unrepresented languages may stem from cross-lingual transfer from other languages. Qwen’s technical report only mentions the inclusion of en and zh in its pre-training data. We conjecture the inclusion of 10 other languages (fr,es,pt,vi,ca,id,ja,ko,fi,hu) based on its significant performance in these languages. and even the smallest models surpass $\tau=65\%$ accuracy in them. These results confirm that LLMs effectively encode value concepts in a multilingual context. Moreover, Figure 1 shows a clear pattern that increasing model parameters and language resources notably improves concept recognition accuracy. This indicates that both expanding the model size and increasing the scale of language pre-training data contribute to a more precise encoding of value concepts.
Appendix E.1 compares the PCA-based method with the mean-based method outlined in §3.1. It reveals that both methods produce concept vectors of comparable precision, with the mean-based technique holding a slight edge. The consistent performance across various extraction techniques confirm the effectiveness of concept vectors in capturing conceptual information. Appendix E.2 indicates that middle layers are particularly adept at capturing abstract concept information. Appendix E.3 demonstrates that even a small number of training samples can effectively extract representations of value concepts in LLMs. For detailed results on each value concept and additional discussions, please refer to Appendix E.4 and E.5.
@@
Figure 2: Cross-lingual similarity of concept vectors across all language pairs, averaged over all value concepts. The languages included in each model’s pre-training data are presented and sorted based on their proportions in the corresponding model’s pre-training data. For Qwen-chat series, we conjecture its language inclusion based on multilingual concept recognition accuracy (§4.2) and display its primary languages, zh and en, at the forefront.
@@
4.3 Q2 & Q3: How Consistent and Transferable are Value Concepts across Languages, and What is the Impact of LLMs’ Multilinguality?
@@
Through computing cross-lingual similarity of concept vectors (§3.3) and recognizing cross-lingual concepts (§3.4), we investigated the cross-lingual consistency and transferability of these value concepts (Q2). Moreover, analyzing these concepts on LLMs trained with different multilingual data distributions provides insights into the multilinguality of LLMs (Q3).
@@ 4.3.1 Trait 1: Inconsistency of Concept Representations between High- and Low-Resource Languages
Figure 2 illustrates the cross-lingual similarity of concept vectors captured by the three 7B-sized models. We find that different multilinguality leads to different patterns of cross-lingual concept consistency. In the case of LLaMA2-chat-7B, the absolute dominance of English results in the model learning relatively independent concept representations for English, showing concept representation inconsistency between English and other languages, while higher cross-lingual concept consistency is observed among other languages. BLOOMZ-7B1’s cross-lingual concept consistency exhibits a very different pattern: the four languages with the lowest proportions (ta, te, sw, ny, accounting for 0.50%, 0.19%, 0.015%, and 0.00007% of pre-training data, respectively) show the lowest concept consistency (similarity) with other languages, while languages with relatively higher proportions (en with the highest percentage of 30.04%, and ca with the lowest percentage of 1.10%) demonstrate higher concept consistency with each other. [^3] We observe inconsistency between Spanish and other languages in BLOOMZ-7B1. We would like to explore this in our future work. For Qwen-chat-7B, we do not observe significant cross-lingual consistency between the main languages (zh, en) and other languages. In summary, cross-lingual concept inconsistency is more likely to occur between high- and low-resource languages.
The findings from Steck et al. (2024) suggest that a high average cosine similarity might raise concerns when dealing with unrelated representations. However, the results in Appendix G.1 indicate that, in our specific context, cosine similarity between concept vectors could reflect their genuine correlation. Additionally, Appendix G.2 presents cross-lingual consistency across different model layers, revealing that intermediate model layers exhibit greater consistency. For comprehensive results on each value concept and further discussions, please refer to Appendix G.3 and G.4.
@@
D. C. D. C. D. C. D. C.
@@
Table 1: Pearson correlation between cross-lingual concept consistency and linguistic similarity for all language pairs. Scores greater than or equal to 0.2 are highlighted in bold. “D.” refers to results obtained through direct computation; “C.” pertains to the average results derived by first categorizing languages based on language resources and then computing correlations within different language categories.
Figure 3: Cross-lingual concept transferability across all language pairs, averaged over all value concepts. Languages are sorted based on their percentages in the pre-training data.
@@ 4.3.2 Trait 2: Linguistic Relationships Distortion due to the Imbalance of Language Data
We employed two calculation methods to estimate the correlation. The first method directly computes the Pearson correlation on all language pairs (Direct), while the second starts by categorizing language pairs based on language resources. Subsequently, correlations are computed within different categories and averaged (Category). Please refer to Appendix F for details of the latter method.
@@
In terms of linguistic variations, cross-lingual concept consistency exhibits the strongest correlation with genetic and syntactic similarity. In contrast, there is a weak positive correlation between cross-lingual concept consistency with geographic similarity, while no correlation is observed with phonological similarity. The results suggest that LLMs embed more consistent value concepts for language pairs with similar syntactic structures, genetic relations, and geographic proximity, aligning with previous findings on multilingual factual knowledge (Qi et al., 2023).
@@ 4.3.3 Trait 3: Unidirectional Concept Transfer from High- to Low-Resource Languages
For a given source language $l_{1}$ and target language $l_{2}$, we compute $\textnormal{Acc}_{c}^{l_{1}\rightarrow l_{2}}-\textnormal{Acc}_{c}^{l_{2}}$ (the difference in accuracy scores) to measure the transferability of concept $c$ from $l_{1}$ to $l_{2}$3.4). We average differences in accuracy scores over all value concepts to measure the overall transferability. If the average difference is greater than 0, it indicates positive transferability from $l_{1}$ to $l_{2}$.
We present the cross-lingual concept transferability of the three 7B-sized models in Figure 3. It provides insights into the influence of LLMs’ multilinguality. Firstly, based on the results of LLaMA- and Qwen-chat-7B, we observe a pattern of monotonic concept transfer from the dominant languages to other languages. This pattern also exhibits an upper triangular cross-lingual transferability (the dashed triangular in Figure 3), indicating that cross-lingual concept transfer from high- to low-resource languages is more prevalent. In contrast, BLOOMZ-7B1 exhibits a relatively balanced bidirectional cross-lingual concept transferability, while for languages with extremely low resources, the tendency of unidirectional transfer persists.
While evaluating transferability based solely on changes in accuracy may introduce biases due to initial performance variations across languages, potentially amplifying the observed unidirectional transfer, Appendix H.1 indicates that transferability is not solely determined by language performance. For comprehensive results on each value concept and further discussions, please refer to Appendix H.2 and H.3.
@@
LLaMA2 -chat-7B No-Control 0.97 1.94 6.80 1.94 6.80 4.85 8.74 5.83 3.88 10.68 14.56 4.85 6.44
@@
LLaMA2 -chat-13B No-Control 0.97 0.97 5.83 1.94 5.83 5.83 27.18 8.74 2.91 10.68 15.53 6.80 8.38
@@
En-Control 88.35 99.03 95.15 98.06 97.09 98.06 93.20 94.17 99.03 97.09 90.29 87.38 95.32
LLaMA2 -chat-70B No-Control 0.00 1.94 4.85 0.97 6.80 2.91 27.18 11.65 2.91 20.39 18.45 10.68 9.89
@@
5 Q4: Is Value Alignment of LLMs Controllable across Languages?
LLaMA2-chat models, trained with alignment techniques such as RLHF, exhibit value alignment capabilities like rejecting harmful instructions. In this section, we employed the Representation Engineering (RepE) methodology (Zou et al., 2023a) to bypass such defense and further explored the potential for cross-lingual control of value alignment.
@@ 5.1 Cross-Lingual Value Alignment Control
To control a LLM to exhibit behavior aligned with a value concept $c$, a straightforward RepE-style method is multiplying the previously extracted concept vector $\bm{v}_{c}$ by a control strength $s$ and adding it to the hidden states of multiple layers $L$ within the target model. This procedure is iteratively applied to each token, formulated as $\bm{h_{i}}^{{}^{\prime}}=\bm{h_{i}}+s\cdot\bm{v}_{c}$, where $\bm{h_{i}}$ and $\bm{h_{i}}^{{}^{\prime}}$ denote the original and perturbed hidden state of $i$-th token, respectively. [^5] Reflecting on §3.1, each layer has its specific concept vector, and the perturbation is executed across multiple layers $L$. We omit the detail here for simplicity. In a cross-lingual scenario, we leverage the concept vector $\bm{v}_{c}^{l}$ of the source language ${l}$ to control the model’s behavior across various target languages. To determine appropriate control strength $s$ and control layers $L$ for cross-lingual control, we first conduct hyperparameter search to choose the combination that demonstrates the most effective control on language ${l}$. Subsequently, we employ this combination for cross-lingual control across all target languages and evaluate the control effect on each of them.
In our experiments, a successful control is steering the LLM to follow a harmful instruction rather than rejecting it. We compute the Following rate, representing the proportion of harmful instructions the model follows, to assess the effectiveness of model control. Specifically, we utilize the multilingual negative testing data (harmful instructions) for the concept of harmfulness (§4.1), calculating the Following rate in each language. Please refer to Appendix I for details of hyperparameter search and model control evaluation.